Monday, July 23, 2012

Day 57, 7/23/2012 -- BEAMS

Last night was great success for overnight team. Make many beams, is good quality!

Ahem...yeah, so apparently last night went pretty well! This morning I came in and climbed the water tower to bring the source down while Franky did some noise tests. I ended up accidentally unplugging some of the lighthouse lights, so I had to climb the water tower again in the afternoon to fix that (loose cables). Franky warmed the focal plane to recover the wafers for one of the crates that overheated yesterday, and it is now cooling so as to be ready for more scans tonight. The agenda tonight is to the first attempt at our absolute polarization rotation measurement.

In comings and goings, this morning Amber and Britt headed back to New York while Michele is being picked up by KyleH at the airport. And, sadly (for you!), I'm taking a break from Palestine for a little while, so the updates will pause here until I return.

I don't have any EBEX pictures today, but I took a couple more BLAST and SuperTIGER pictures, which you can see in their respective albums:

Sunday, July 22, 2012

Day 56, 7/22/2012 -- Sunday sun day

Various things done today -- Britt and I got the two solar arrays fully populated with solar panels, we cycled the fridges, we got the bolo system connected to its flight power system with a battery (built-in UPS!), filled helium, and the night shift is getting ready to take more water tower scans overnight tonight.

In the morning, I had to climb the water tower to retrieve a piece of equipment that I forgot to bring down yesterday; it had to weather the storm (covered by a plastic box) and I had to make sure it still worked. And, despite my bumbling, it did, even after I accidentally connected the +12V power to the output and tried to measure the output signal from the 12V input power connector. Oops.

The overnight team had a dramatic victory last night, though. As I mentioned in yesterday's post, there were thunderstorms in Palestine overnight. The power at the high bay often goes out during these thunderstorms, and if the bolometer system loses power, we lose all of our bolometer biases, the bolometers latch superconducting, and the detectors have to be warmed en masse in order to recover them, a process that takes many hours as the focal planes cool again to their operating temperature. Concerned about this possibility, Kyle insisted that they hook the bolometer system up to a set of batteries in order to prevent this in case the power did go out, as is its wont. After frantically assembling a set of cables and going through the switchover procedure, they got the system on batteries -- and not 2 minutes later, the high bay power went out. But the system stayed powered!


Saturday, July 21, 2012

Day 55, 7/21/2012 -- High and low

In order to manage gondola time more effectively, Franky and I are shifting to an 'early' schedule where we get in to the high bay around 7AM (rather than 9AM as we had been previously). This allows Franky more time to mess with detectors after the end of the night's scanning at ~6AM, and means I get to go up on the tower and bring down our fragile and expensive millimeter-wave source before the temperature starts rising.

So after an invigorating climb up the water tower, I set about doing some gondola work -- I managed to install the set of brackets for the second power box and the battery tables underneath the gondola, necessitating a lot of crawling around on the floor and drilling holes. Once Jeff came in in the afternoon, we also took a set of mirror position measurements to see if they are actually where they should be after having aligned the mirror mounts.

Franky was able to recover a fair number more detector combs, so we have nearly our full complement of detectors available for testing. There are thunderstorms tonight, so there likely won't be any scanning tonight. However, the night crew has some other tests they can do, so it's not a wasted night.


Friday, July 20, 2012

Day 54, 7/20/2012 -- SIPpin' on gin and juice

Today our Science Instrumentation Package, or SIP, was moved into the high bay to prepare for integration with EBEX. The SIP is our interface to the experiment, handing things like command uplink and data downlink (over a variety of different protocols: line-of-sight biphase, TDRSS satellite relay, and Iridium satellite) as well as a small number of high-reliability hardware on/off comands and all of CSBF's payload telemetry. As this is a new SIP, there will be a bit of testing first before it gets installed on the gondola.

After Franky finished investigating some SQUID issues in the morning, we cycled the fridges again (yesterday's cycle didn't last long due to the need to keep the temperature warm for a long period of time for debugging). I made a cable to connect a set of ground batteries to the EBEX bolo power system so we can run the detectors off of cleaner power than the switching power supply that we currently use. I also did a little bit of work on the power system filter box, but found I was missing some connectors Linkthat I need in order to put it together for reals (since ordered!). Britt worked on more parts for the flight power system as well.

The plan for the evening is to take a first cut at doing beam maps. The source will be mounted up on the tower tonight, and the night crew will try and find a detector so they can set levels (to avoid saturation) and hopefully do some first beam maps.


Thursday, July 19, 2012

Day 53, 7/19/2012 -- There! Are! SIX! Lights!

Today we ran the first fridge cycle with the cryostat on the gondola. Things mostly went well (it's pretty awesome to send a single command and have the gondola tilt to the right position and then run the cycle completely automatically) except for a few issues biasing the detectors towards the end. Franky spent the afternoon debugging what went wrong with the detector biasing.

The other big ticket item today was setting up the so-called 'lighthouse' on the water tower, requiring the first of what will be many (many) trips up to the top of the tower. Jeff had the clever idea of using small power bricks to power the lights this year, so instead of lugging two ~25 lb power supplies up the 150 meters of the water tower, we only had to take six ~1 lb (or less) 12VDC power bricks up (along with the rest of the light hardware). KyleH and I got the whole thing set up in about 2 hours, so now we have six lights up on the water tower to use as a 'fake sky' for pointing while doing water tower scans.

On the agenda for tonight is a relative timing test between the ACS and bolo systems as well as scans of the lighthouse to build a source catalog to use in subsequent scans.

Picture (just one -- I'll take more pictures from atop the tower when I don't have to lug other equipment up too):

Wednesday, July 18, 2012

Day 52, 7/18/2012 -- Water tower scans soon

We are putting the final pieces in place before we are able to start taking calibration measurements with the integrated system.

Jeff and I took a look at the millimeter wave source we use to do a variety of calibration measurements; we were able to couple it to a function generator in order to do time detector constant tests over the full range of frequencies we want to measure. After the time constant tests, the source will then be placed atop the CSBF water tower and aimed down at the high bay and we'll start doing beam maps and our absolute polarization angle calibration. To that end, I also worked out the coordination with Verizon Wireless that they would turn down their cell phone transmitter at our request for the next 3 weeks in the late night/early morning so we can take data without seeing the interference from the tower.

I also managed to route the final dryer hose (containing the cryostat housekeeping signals) into position, and Kyle finished the securing and cable connections. Kyle and Britt also worked all day on checking the system grounding and ensuring the bolo and ACS systems weren't inadvertently shorted together. Britt also did a lot of work on the power system to get it ready to run off batteries when we need to.

Franky spent the day working with the SQUIDs -- the first detector system work that's been done since moving the cryostat to the gondola.


Tuesday, July 17, 2012

Day 51, 7/17/2012 -- Mirror, mirror, on the...gondola?

Today was another insanely productive day. Jeff and I aligned the primary mirror mount in the morning, Franky and others were able to mount all of the detector readout cabling in their metallic 'dryer hoses' (which currently look pretty terrible since we didn't see the point in mangling fresh ones and having to replace them again in Antarctica), Jeff, Michele, Kyle, KyleH, Britt, and I mounted both mirrors, Kyle routed and secured the dryer hoses so they wouldn't interfere when the gondola tilts down in elevation, and Jeff and I installed and measured the position of an inclinometer atop the cryostat to be used in the polarization rotation calibration test. And then Amber's husband grilled dinner for everyone!

Not bad for a day's work.

Cynthia left Palestine today to return to MN to teach a summer course, so we're a little shorthanded at the moment.


Monday, July 16, 2012

Day 50, 7/16/2012 -- Alignment, test fitting, network debugging

Today we had progress on multiple fronts simultaneously. Jeff and I spent a fair bit of today on the gondola atop the cryostat making measurements and adjustments for our mirror mounts. We got the secondary mirror mount aligned and will tackle the primary mirror tomorrow.

Elsewhere in the collaboration, Andrei and KyleH continued to set up hardware for the Bemco test of the liquid cooling system. Kyle plugged in the final set of BRO cables (the gondola network optical fibers), and he and Seth spent the rest of the day debugging network communications issues (successfully).

And last, the Baffle/Triangle Support (BTS) was test-fitted on the gondola along with the triangle spreader bar we use and we verified that everything fits properly (only minor 'convincing' was required to get all of the holes to line up).

I've posted pictures from yesterday in yesterday's post if you want to go back and look at them.

Pictures from today:

Sunday, July 15, 2012

Day 49, 7/15/2012 -- Crises, crates, and cables

I realize that I forgot to mention our little liquid helium crisis back on Friday, so I'll mention it here, because the crisis has been summarily averted:

On Friday, we needed to fill the liquid helium tank of the EBEX cryostat. This was no problem, as we had received a full 500-liter dewar about a week and a half prior that would be sufficient to fill EBEX nearly four times over. The only problem wasn't full. In fact, it wasn't anywhere near full. Actually, it turns out it was completely empty. And warm. We found this out Friday morning, and the helium in EBEX would run out on Sunday if we didn't get our fix. And if the helium ran out, the cryostat would start to warm up...and that's not good at all.

So...we let CSBF know, and they set the wheels in motion to see if we could get helium over the weekend. From their liquid helium east Texas. They ended up having two CSBF guys drive the 11-12 hours out to the vendor where a full 500L dewar would await them at noon on Saturday, and then drive all the way back so we would have our helium by midnight Saturday. they did! We were able to fill our helium before the situation got really dire -- CSBF really goes out of their way to make sure we're able to get science done.

Aside from helium, today was largely a continuation of yesterday's work. We started adding the cables for the bolo system to the gondola in order that they could be routed in a usable way without risk of them tearing or being sheared by gondola motion. While doing this, I noticed one of the bolo system power cables was noticeably lighter than the others -- it turns out it had been made with the wrong gauge wire (too small), so I set Cynthia to work on making a new cable while Jeff and I worked out the issues routing the other cables. Jeff also mounted the HWP control and readout crate. Kyle spent most of the day debugging commanding issues in Power Crate 1 (with remote help from Ilan over the phone) and seems to have worked out all of the known problems. So with that done, all of the major components of the bolo system were in place on the gondola with the exception of the readout cables.

Elsewhere in the high bay, Andrei (oh, forgot to mention, he was gone and came back on Thursday) and KyleH started setting up for a Bemco test of our liquid cooling system components. In other comings and goings, Franky arrived from McGill late last night to replace Kevin, who will be leaving tomorrow morning. And Kate and daughter left around mid-day today.

I have pictures, but I forgot to grab the memory card from my camera so their posting will have to wait until tomorrow.

UPDATE -- Pictures:

Saturday, July 14, 2012

Day 48, 7/14/2012 -- I gotta say, today was a good day

Some days, everything just goes right. We came in early, broke down the detector readout system, installed the double window parts, removed the cryo mass dummy from the gondola, and installed the cryostat into the gondola -- all in 3 hours. The rest of the day was devoted to mounting the detector readout boxes (the "BROs"), of which we got all 4 mounted, mounting one of the power crates (the other needs some investigations), and tying up some of the cables so Joy and Chappy can do scan tests tonight without having to worry about snagging cables.

All in all, pretty awesome.

And turns out we're not the only ones who have a cryostat mounted on their gondola -- the BLAST team mounted theirs yesterday.


Friday, July 13, 2012

Day 47, 7/13/2012 -- The end is near!

The end of this phase of integration, that is -- Kyle is completing the final Ebert-Fastie scan tonight; I did one scan earlier and completed our suite of polarization modulation efficiency measurements.

Jeff returned from Minnesota today as well; his first task was to look at some of the data we took in his absence. Thankfully, it appears to be usable, so we don't have to re-do it! Jeff also put Cynthia to work making some cables for our upcoming absolute polarization angle calibration measurement.

The gondola crew started making preparations for the upcoming 'big lift' tomorrow -- putting the cryostat on the gondola. Tomorrow morning we head in early to start disconnecting cables and get the cryostat ready to mount on the gondola. So that means it's time for me to go to bed!

No pictures today, as I was too busy taking data to take pictures.

Thursday, July 12, 2012

Day 46, 7/12/2012 -- Burning the candle at both ends

It has been a long week here in the high bay for Kyle and myself -- but more so for Kyle. He has been on the late shift, getting in at about 3PM and working until 4, 5, or 6AM taking scans, babysitting fridge cycles, and analyzing data. It has taken its toll, and we're both looking forward to being done with this phase of the experiment.

Today was, of course, more scans, one of which seemed to be a waste of time (low signal-to-noise), the other quite reasonable. We tried to start taking another scan, but in the interests of Kyle's health and sanity, we are fridge cycling 'early' (which means Kyle will be in bed by 3AM!) and we will take the last of this calibration data tomorrow. After that, we break down the instrument (move the Ebert-Fastie, disconnect cables, etc) in preparation for the cryostat going on the gondola Saturday morning.

Britt and Michele finished Bemco'ing the power system, which passed with no problems (as expected based on the specs of the components). Kevin put Cynthia to work fleshing out our focal plane visualization tool, which will be nice to have soon! I also started work on a filter box to clean up the power coming out of the bolo system DC/DCs.


Wednesday, July 11, 2012

Day 45, 7/11/2012 -- Nearing full integration

Today I managed to take some crucial data, that of our polarization modulation efficiency -- how well does our receiver act as a polarimeter. Pretty darn well, it turns out! But I spent too much time in the morning trying to get some spectral data on one of our 410GHz detectors without much luck -- hopefully Kyle will have success doing scans during the late shift.

Britt and Michele successfully tested our elevation actuator and our inner frame lock pin actuator in the Bemco. The power system (charge controllers, batteries, and power distribution boxes) are on the agenda for tomorrow.

Tomorrow will also likely be the final day of cryostat-only tests -- we're hoping to have the cryostat put on the gondola by the end of the week!

No pictures today (I was too busy!).

Tuesday, July 10, 2012

Day 44, 7/10/2012 -- Groovin'

It's amazing how much data one can take when things actually work and you can get in the groove. We did more scans of another wafer today, and when one gets massive signals it's easy (and quick) to take data. We did 4 or 5 separate scans today, both with and without the half-wave plate rotating. During the scans, Kyle worked more on his analysis code (after he arrived at the high bay in the afternoon), I helped Kevin fix some DfMUX boards and helped Shaul analyze some data taken back in MN, and Kate and Kyle continued working on understanding our bolometers.

On the gondola side, Britt and Michele started setting up hardware in the Bemco for a test of the new power system's electronics and batteries. In addition, they are going to test the inner frame lock pin and the spare elevation actuator, and BLAST is testing a shutter for their cryostat (apparently there are times during their flight where they expect to be pointed toward the sun...and the sun focused through a 1.8-meter mirror generates a pretty high power density which would easily melt the window of their cryostat). On the late shift, Chappy and Joy will be working on testing scan modes while Kyle babysits the fridge cycle to ensure we have working cold detectors for tomorrow's tests.

Last, I took a peek around the last experiment that will be flying from Antarctica this year, SuperTIGER. If you're too lazy to click on the link, it's a heavy-ion cosmic ray detector with an aim to help constrain the possible sources of high energy cosmic rays by measuring the ratios of the various elements that make up cosmic rays. I took some pictures, and like the BLAST album, it will be updated as necessary but probably won't be mentioned on this blog much:


Monday, July 9, 2012

Day 43, 7/9/2012 -- Stormy weather

Last night, Kyle stayed in the high bay overnight to make sure the fridge cycle went off without any issues -- and it's a good thing too, because due to some files not having been updated correctly, the detectors wouldn't have been tuned. Because of Kyle's vigilance, he was able to tune the detectors by hand and put us in a good position this morning to take some measurements.

The goal for today was to take some more spectral measurements at 250 and 410GHz. I started at 410, since that's how the spectrometer was set up already, but was having no luck finding the signal. It was only later, after switching the hardware for 250 GHz measurements, that I had the grating in completely the wrong position, so it's no wonder I didn't see anything. Still, the 250GHz measurements were on track for going much more smoothly (we found a detector with a whopping big signal)...and then I broke the motor for the chopper wheel. After a frantic bit of hunting around for the replacement motors, we finally were able to take some data.

We had a pretty massive downpour today, which caused the power to go out -- but luckily, we now have the whole bolo system running off of a UPS, so we didn't lose detector biases and were able to continue with measurements without much delay. Awesome!

On the logistical front, today we had a long meeting trying to figure out our schedule for the rest of the integration -- mostly, how long do we keep doing these spectral measurements before putting the cryostat on the gondola. Right now it looks like we have just a couple more days to wrap up these measurements.

On a blogistical note, it's getting harder for me to know of who all is doing what around the high bay, especially as there are few visible changes (the gondola's mostly put together and the receiver looks the same as always), and people are now working different schedules to try and make the most efficient use of time. Thus it's hard to keep track of what's going on without wandering around and asking people -- which is tough when I've got work to do! So of course all of my updates focus on receiver stuff since that's what I'm working most directly on. C'est la vie.

Finally, we had another collaboration member join us -- Amber, the EBEX Co-I from Columbia, arrived with her newborn baby and husband in tow.


Sunday, July 8, 2012

Day 42, 7/8/2012 -- A wretched hive of scum and villainy

The EBEX bolo system power crates, that is -- designed years ago, they are probably the worst-laid-out piece of major equipment on the experiment. One of them is, anyway -- power crate 1, which houses DC/DC converters for half of the bolo system, the housekeeping system, the bolo side network, and a general housekeeping board which, among other things, controls various parts of the HWP system). Needless to say, this crate is very crowded, which means that we try as hard as we possibly can to not mess with it, lest we be forced to delve into the bowels of the crate.

Well, we were forced to. In order to try and solve a nagging DfMUX board power issue, we had to swap out the DC/DC voltage setting resistors for the +/- 5.8V power system, which entailed opening up the crate (as much as we could) and swapping out these resistors to drop the voltage by about 200mV. Along the way, we found that a power return wire for one of the DC/DCs was no longer connected and was subsequently fixed.

Power crate fixes, though, happened at the end of the day (we are just wrapping up now, at midnight). During the day, we were able to take spectral measurements for two 410GHz detectors. Kate and Kevin are now on the 'early' shift (7AM - ~4PM), so they could get some time with the DfMUX system without me and Kyle trying to get data.

The gondola crew spent much of today doing gyro tests, calibrating each individual gyro (there are six total, two full three-axis redundant gyro boxes) by swapping each gyro to a position where it would sense elevation motion and doing careful elevation slews. Britt worked on a little electronics board for the solar power system, and Michele worked on straightening out a small ding in our spare actuator's outer housing tube.


Saturday, July 7, 2012

Day 41, 7/7/2012 -- Dealing with things not working

We came in today to warm cold stages -- this time, neither of our fridge cycle commands worked overnight. After debugging in the morning by Seth, we started new fridge cycles around noon. Most of Kyle's day was spent analyzing old data while Kevin and I worked on debugging issues in the readout electronics. Kyle, Cynthia, and I are back in the high bay at ~9:30PM to try and start a scan and hopefully get some data tonight.

The gondola had its legs attached and the team is doing some initial shakedown tests.

In addition, Kate returned to Palestine, this time with her 1-year-old daughter in tow. Kate and Kevin plan on working out detector issues in the ~week that Kate will be down here.

Pictures (not many):

Friday, July 6, 2012

Day 40, 7/6/2012 -- Gondola the White returns

Today's big event was the joining of the inner and outer frames and the mounting of our cryostat mass replacement, the so-called cryodummy. Basically, because of all the work the gondola crew did on mounting electronics and other hardware when the two were apart, now that everything is back together they are ready to hang and start doing scan tests!

On the receiver side, we had a partially botched fridge cycle (our 1K fridge didn't cycle at all), so that limited the amount of data we could take. Still, we managed to get some. On the downside, one of our DfMUX boards let out some of its magic smoke (from a capacitor). After a little investigation, the cause was found and it will be fixed forthwith. The fridges shall be cycled again tonight (with the issue that prevented last night's cycle found and corrected), so we should be good to go again tomorrow.


Thursday, July 5, 2012

Day 39, 7/5/2012 -- Everything's coming up Milhouse

After the past few days of battling with electronics and fridge issues, it was a refreshing and welcome change to come in to the high bay this morning with everything working. We are taking more polarized Ebert-Fastie data, and indications so far are that we can use data taken during the polarized scans to get both polarization data and overall power spectral response information, which will make our data gathering twice as efficient!

Kyle returned from Minnesota late yesterday, bringing with him a new UMN grad student, Cynthia, who's here to help out where possible and gain experience in the field. One of the first things Kyle noticed when arriving at the high bay was that the mirrors had been put on the inner frame -- at which point he pointed out that we needed to align the mirror mounts first. And aligning the mirror mounts requires having the cryostat on the gondola, as the mirror mounts (six-legged contraptions known as "hexapods" or "Stewart platforms") are aligned relative to the optics inside the cryostat. So, off they came.

But the gondola team got plenty of other stuff done today: They mounted the trunnion bearings, mounted the flight computer and ACS crates, and wrapped the flight ropes (those sexy lavender beasts) with red balloon wrap to protect against abrasion and other damage.


Wednesday, July 4, 2012

Day 38, 7/4/2012 -- Fourth of July

Today we spent time dealing with bolo electronics -- an unexpected reboot of the crates overnight meant that the overnight tuning didn't work, and later we had some other board issues that we've decided to debug more fully before proceeding further. That took up most of the day for myself, Jeff, and Kevin. In addition, Jeff has spent a lot of time refining his data analysis of our polarized Ebert-Fastie data.

The gondola team has made great progress today, mounting both mirrors and the primary mirror protection rollbar along with a few other ancillary pieces of gondola hardware (gyro box, star camera mounts, elevation actuator, etc). Tomorrow they plan on setting the inner frame atop the outer frame, joining the two major pieces of the gondola together again.

In other news, I've decided to move all BLAST-Pol pictures into a separate web album to avoid confusion with EBEX hardware. I will continue to update it but I probably won't post the link here again unless there's something particularly notable. You can see that album here:

Today's (EBEX) pictures:

Tuesday, July 3, 2012

Day 37, 7/3/2012 -- Not awesome

Today was kind of a bust in terms of getting any real data. Last night after dinner, Seth and I went back to the high bay to do a simple fix on one of the "general housekeeping" boards in the system that lives in one of the bolometer power crates. Turns out that after we reconnected everything, the cryostat temperature readouts stopped working. And I was informed of this at approximately 1AM. So I went back to the high bay to try and fix it, was thoroughly unsuccessful, Seth tried to debug the issue in the morning before we got in, and we finally tracked it down to connecting mislableled connectors on the power box the way they were indicated -- i.e. incorrectly!

After fixing the labels, we then set in motion a new 'schedule file' for the bolometer system that would automatically put the detector array into a nearly-operational state. Of course, this didn't work completely initially, leading to some time spent debugging. And then the focal plane was taking forever to cool...partially because one of the fridges ran out. Even so, we pushed ahead and tried to get some data as the detectors were still quite cold...but we ended up not being able to get any usable calibration data today. Major bummer. On the plus side, Jeff was able to make some progress on the data analysis for our calibration scans giving us some more confidence that our data is of good quality.

The gondola team continues to make progress; now that the Bemco testing is done they have started to reassemble the entire gondola. A continuing discussion with CSBF revolves around placement and protection of their SIP LDB electronics package, and I believe they are converging on a solution.


Monday, July 2, 2012

Day 36, 7/2/2012 -- Bemco day!

Today's big event (for us) was the ACS team testing a whole bunch of stuff in CSBF's Bemco environmental test chamer that replicates the conditions during various phases of the flight. Despite a slow start, they managed to get all of their testing done in one day, which means they can now start putting the gondola back together.

On the receiver side, Jeff and I continued to do polarized Ebert-Fastie scans. We need statistics, sue us.

The other big news around the base is that one of the other payloads, BLAST-Pol, showed up. One of the early BLAST (pre -Pol) blogs was my inspiration for doing an EBEX blog, and now they're just across the building from us in the west high bay (sadly, where the badminton court was). In addition, EBEX's gondola design is in some ways derived from BLAST's, so don't be surprised if you see similarities between the two in the pictures.


Sunday, July 1, 2012

Day 35, 7/1/2012 -- Couldn't think of a title!

Today started off not particularly well -- we had set the fridges to cycle overnight so we wouldn't be caught with a fridge that ran out in the middle of taking data. This all worked perfectly fine, except we noticed that one of the fridges started warming up immediately after we got in to the high bay in the morning. It was then that we realized we forgot a crucial step in the cycling process. Because the cryostat will, on average, live tilted 25 degrees off vertical during flight, the fridges inside are mounted so that they will be vertical when the cryostat is tilted 25 degrees. When the cryostat is NOT tilted, the fridge cycle...doesn't really work (due to the geometry of the particular fridges we use). And, well, we forgot to tilt the cryostat last night before we left.

So we ended up running the fridge cycle again this morning, at 9AM after tilting the cryostat. And the fridge cycle takes about 5.5 hours just to run...and then a few more hours before the focal planes are stable at their base temperatures, which is necessary before we can start working with the detectors. Which basically meant that for me and Jeff, our day was completely shot in terms of taking receiver calibration data. However, we did manage to make some parts to upgrade the Ebert-Fastie: we use a servo-controlled mechanical chopper wheel to modulate the signal coming out of the Ebert-Fastie. However, we noticed that the chop frequency wasn't always very stable, so Jeff had the idea to try coupling the chopper wheel to a stepper motor rather than a servomotor. This ended up being surprisingly easy, since Jeff had on hand some small stepper motors that had the same shaft size as the servomotor, and a spare stepper motor controller. All that needed to be done was to wire up the controller (Jeff's task) and make a new motor mount (my task). And, lo and behold, it all works! It looks like the chopper is MUCH more stable with a stepper motor driving it.

Elsewhere in the high bay, the ACS team continues to work on getting things ready to Bemco test. They have now reassembled the star cameras and disk pressure vessels and purged them with dry nitrogen before sealing them. And Andrei and KyleH did initial tests of wrapping the BTS with its aluminized mylar skin.

In the next few days, the other payloads headed to Antarctica will be showing up, so we had to move all of our miscellaneous stuff out of the west high bay and find a home for it in our high bay.